OpenStudy (anonymous):

Abel uses a probability simulator to roll a six-sided number cube 100 times and to flip a coin 100 times. The results of the experiment are shown below: Number on the Cube Number of Times Rolled 1 12 2 18 3 30 4 22 5 10 6 8 Heads Tails 34 66 Using Abel's simulation, what is the probability of rolling a 2 on the number cube and the coin landing on tails? fraction 84 over 10,000 fraction 1,188 over 10,000 fraction 18 over 100 fraction 66 over 100

OpenStudy (anonymous):

can i get help by help step

OpenStudy (anonymous):

@Mertsj can you help

OpenStudy (mertsj):

Probability of A and B for independent events is P(A) x P(B)

OpenStudy (anonymous):

36*66?

OpenStudy (anonymous):

2376/2 = 1188? @Mertsj

OpenStudy (mertsj):

What is the probability of a 2?

OpenStudy (anonymous):

100/2?

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

am i on the right track

OpenStudy (anonymous):

@kropot72 am i on the right

OpenStudy (kropot72):

The probability of rolling a 2 is given by: \[\large P(2)=\frac{number\ of\ 2s\ rolled}{total\ rolls}\]

OpenStudy (anonymous):

ok @kropot72

OpenStudy (anonymous):

1188/10000 @kropot72

OpenStudy (kropot72):

@steffie84 How did you get that?

OpenStudy (anonymous):

66*34/2*100*100

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Not really. What is the probability of rolling a 2 (using the equation in my first post here)?

OpenStudy (anonymous):

total roll is 100?

OpenStudy (anonymous):

@kropot72 am not sure what

OpenStudy (kropot72):

The number 2 came up 18 times in a total of 100 rolls.

OpenStudy (anonymous):

@kropot72 just saw what you are talking about

OpenStudy (anonymous):

18/100

OpenStudy (kropot72):

Correct. The probability of tails is found from: \[\large P(tails)=\frac{number\ of\ tails}{total\ flips}\]

OpenStudy (anonymous):

66/100

OpenStudy (kropot72):

Correct again. Therefore as @Mertsj told you, the required probability is given by: \[\frac{18}{100}\times\frac{66}{100}=you\ can\ calculate\] Hint: Do not simplify the result of your calculation.

OpenStudy (anonymous):

1188

OpenStudy (anonymous):

10000

OpenStudy (kropot72):

So the result is \[\large \frac{1188}{10000}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@kropot72 @Mertsj thanks

OpenStudy (kropot72):

You're welcome :)